### v1

# Text2Cypher Across Languages: Evaluating Foundational Models Beyond English

Abstract.
Recent advances in large language models have enabled natural language interfaces that translate user questions into database queries, such as Text2SQL, Text2SPARQL, and Text2Cypher. While these interfaces enhance database accessibility, most research today focuses solely on English, with limited evaluation in other languages. This paper investigates the performance of foundational LLMs on the Text2Cypher task across multiple languages. We create and release a multilingual test set by translating English questions into Spanish and Turkish while preserving the original Cypher queries, enabling fair cross-lingual comparison. We evaluate multiple foundational models using standardized prompts and metrics. Our results show a consistent performance pattern: highest on English, then Spanish, and lowest on Turkish. We attribute this to differences in training data availability and linguistic characteristics. Additionally, we explore the impact of translating task prompts into Spanish and Turkish. Results show little to no change in evaluation metrics, suggesting prompt translation has minor impact. Our findings highlight the need for more inclusive evaluation and development in multilingual query generation. Future work includes schema localization and fine-tuning across diverse languages.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The translated dataset is available at HuggingFace: https://huggingface.co/datasets/mgoNeo4j/translated_text2cypher24_testset
1. Introduction
Databases provide efficient mechanisms for storing, organizing, and retrieving data. Query languages such as SQL (for relational databases), SPARQL (for RDF graphs), or Cypher (for graph databases) enable users to interact with these systems (Hogan et al., 2021). Advancements in large language models (LLMs) have enabled the translation of natural language questions into database queries (Text2SQL, Text2SPARQL, Text2Cypher), making databases more accessible to non-experts. However, most research in this area has focused on English, with limited attention given to other languages (Jannuzzi et al., 2024; Geng et al., 2024).
Figure 1. The user wants to generate a Cypher query for the question ‘What are the movies of Tom Hanks‘. However, they may express the same question in different languages, such as English (EN), Spanish (ES) or Turkish (TR).
This work focuses on Text2Cypher, converting natural language questions into executable Cypher queries (see Figure 1). For example, a user might want to write a Cypher query for the question ”What are the movies of Tom Hanks?”, which can be expressed in different languages, such as English (EN), Spanish (ES), or Turkish (TR). Regardless of language, a Text2Cypher model should generate the same query, such as ”MATCH (actor:Person {name: ”Tom Hanks”})-[:ACTED_IN]-¿(movie:Movie) RETURN movie.title”. Furthermore, we aim to evaluate multilingual performance of the foundational LLMs on this task. For this purpose, we construct a test set of semantically aligned questions in English, Spanish, and Turkish, all paired with the same Cypher query. This setup enables cross-lingual benchmarking, isolating the impact of input language on model performance. Our main contributions are:
  • We release a multilingual version of the Text2Cypher test set (Ozsoy et al., 2025), translated into Spanish and Turkish to enable cross-lingual evaluation [^1] Dataset: https://huggingface.co/datasets/mgoNeo4j/translated_text2cypher24_testset.. These languages were chosen for their linguistic diversity and resource differences.
  • We evaluate multiple foundational models on this dataset using standardized prompts and metrics, comparing performance across English, Spanish, and Turkish with semantically equivalent questions.
  • We analyze the impact of language on Text2Cypher performance. Our results show that models perform best on English, then Spanish, and lowest in Turkish. We attribute the lower performance in Turkish to its agglutinative structure and inguistic distance from Indo-European languages.
2. Related Work
Large language models (LLMs) have become central to a wide range of natural language processing tasks. While many early LLMs were trained primarily on English data, recently their capabilities are extended to other languages through the development of multilingual models (Nicholas and Bhatia, 2023). Researchers have also begun exploring the internal mechanisms of these models (Zhao et al., 2024; Kargaran et al., 2024; Zhong et al., 2024; Schut et al., 2025) and evaluating their cross-lingual performance in various application domains (Ozsoy, 2024; Jannuzzi et al., 2024). Multilingual models are trained on datasets that include multiple languages, allowing them to learn cross-lingual representations. This enables them to use word associations and grammatical patterns learned from high-resource languages, such as English, and apply them to languages with less available training data. However, because English still dominates most training sets, these models may also transfer linguistic patterns and assumptions specific to English into other language contexts (Nicholas and Bhatia, 2023; Zhao et al., 2024). As a result, LLMs perform significantly better in higher-resource languages and those similar to them compared to lower-resource languages (Nicholas and Bhatia, 2023).
In the context of translating natural language to database query languages, much of the multilingual work has focused on the Text2SQL task (Dou et al., 2023; José and Cozman, 2021). Recent studies have considered languages such as Portuguese (Jannuzzi et al., 2024; Pedroso et al., 2025), Arabic (Almohaimeed et al., 2024), Russian (Bakshandaeva et al., 2022), and Turkish (Kanburoglu and Tek, 2024). In this work, we study the Text2Cypher task, which translates natural language questions into Cypher queries.
3. Text2Cypher Beyond English
3.1. Dataset Preparation
We used the publicly available Text2Cypher dataset (Ozsoy et al., 2025), which includes English questions, database schema, ground-truth Cypher queries, and metadata. In order to evaluate foundational models across languages, we translated the question field of the test split into Spanish and Turkish using an LLM, and compared results across all three languages. English is an extremely high-resource Indo-European language, Spanish is a high-resource Indo-European language, and Turkishis a medium-resource Altaic language with distinct linguistic traits (Joshi et al., 2020; Nicholas and Bhatia, 2023). Based on this, we expect slightly lower performance for Spanish and a larger drop for Turkish due to linguistic differences. We used the following steps for translation:
  • Masking Named Entities and Quotes: To ensure that named entities and quoted text remained unchanged during translation, we first masked them. Named entities were replaced with placeholders indicating their type and index (e.g., LOCATION_0), while quoted strings were replaced with QUOTE_¡index¿ markers. For example, the sentence ”Hello, I work at ’Neo4j’ in London” was masked as ”Hello, I work at QUOTE_0 in LOCATION_0”.
  • Translation Using an LLM: We translated the masked questions using the GPT-4o-mini model, guided by the prompt shown in Table 1.
  • Restoring Masks: After translation, we replaced the placeholders with their original named entities and quoted expressions to reconstruct the final question.
Note that the ground-truth Cypher queries were not translated. These queries contain (i) Cypher-specific syntax (e.g., MATCH, WHERE), (ii) terms from the schema (e.g., Person, ACTED_IN, Movie), and (iii) user-provided literals (e.g., ”Tom Hanks”), all of which are intended to remain in their original form.
Table 1. Translation prompt for {language_name}
Translation instruction prompt
You are a professional translator. Your task is to translate the given text from English to {language_name}. Follow these guidelines:
1. Maintain the original meaning and tone of the text
2. Preserve any placeholders in the format ¡TYPE_NUMBER¿ (e.g., ¡PERSON_0¿, ¡QUOTE_1¿)
3. Keep the translation natural and fluent in {language_name}
4. Maintain proper formatting and punctuation
5. Do not translate or modify any placeholders
6. Do not translate any text within quotation marks
7. Do not translate named entities (proper nouns, names, places, organizations)
8. Do not translate numbers, dates, or measurements
Please provide only the translated text without any explanations or additional context.
3.2. Experimental Setup and Evaluation Metrics
We conduct our experiments on the test split of the Text2Cypher dataset  (Ozsoy et al., 2025). To support multilingual evaluation, we automatically translated the original English questions into Spanish and Turkish using an LLM. During translation, 50 samples had masking related issues and were excluded from evaluation. As a result, each language was evaluated on the remaining 4,783 samples. We use the following foundational models: (i) Gemma-2-9b-it: Released in June 2024. Primarily supports English. (ii) Meta-Llama-3.1-8B-Instruct: Released in July 2024. Primarily supports English and several European languages. (iii) Qwen2.5-7B-Instruct: Released in September 2024. Primarily supports Chinese and English and maintains multilingual support for over 29 languages. We use their Unsloth versions of models, finetuned with Instruct and quantized to 4-bit precision (e.g., unsloth/gemma-2-9b-it-bnb-4bit) for efficiency and ease of use. For the Text2Cypher task, we use the same prompts as in previous work (Ozsoy et al., 2025), presented in Table 2. After generation, an additional post-processing step is used for removing unwanted text, such as the ’cypher:’ suffix. We use the HuggingFace Evaluate library (Evaluate, 2024) to compute evaluation metrics. We employ two evaluation procedures: (i) Translation-based: Compares generated Cypher queries with reference queries based on their textual content. We report the ROUGE-L score for this evaluation. (ii) Execution-based: Executes both the generated and reference queries on the target database and compares their outputs (sorted lexicographically). We report the Exact-Match score for this evaluation.
Table 2. Instructions used for Text2Cypher task
Type Instruction prompt
System
Instruct.
Task: Generate Cypher statement to query a graph database. Instructions: Use only the provided relationship types and properties in the schema. Do not use any other relationship types or properties that are not provided in the schema. Do not include any explanations or apologies in your responses. Do not respond to any questions that might ask anything else than for you to construct a Cypher statement. Do not include any text except the generated Cypher statement.
User
Instruct.
Generate Cypher statement to query a graph database. Use only the provided relationship types and properties in the schema.
[…]
4. Experimental Results
We evaluate foundational LLMs on Text2Cypher across English, Spanish, and Turkish, focusing on question and prompt language.
4.1. Impact of question language
We first examine the impact of question language on foundational LLMs using the translated dataset. In this stage, we do not modify the task prompt for the Text2Cypher task (shown in Table 2), which is kept in English. Figure 2 shows the performance of the foundational models on English, Spanish, and Turkish questions. The results indicate that all models perform best on English questions, followed by Spanish, with the lowest performance observed for Turkish. This pattern aligns with differences in language resourcedness: English is an extremely high-resource, Spanish is a high-resource, and Turkish is a medium-resource language (Nicholas and Bhatia, 2023; Joshi et al., 2020). LLMs generally perform better on languages with richer and more diverse training data. Additionally, prior work (Dhamecha et al., 2021) shows that models generalize more effectively across linguistically similar languages. As Spanish and English both belong to the Indo-European language family, while Turkish belongs to the Altaic family, structural differences may further contribute to performance disparities. In these experiments, the prompts remained in English, which we hypothesize may have introduced confusion for the LLMs when processing non-English questions. In the next section, we explore this by presenting results for one of the foundational models using both translated questions and translated prompts.
(a) Translation-based evaluation: ROUGE-L score
(b) Execution-based evaluation: Exact-Match score
Figure 2. Performance of models on Text2Cypher when input questions are in English (EN), Spanish (ES) or Turkish (TR).
Table 3. Spanish (ES) and Turkish (TR) Instructions used for Text2Cypher task
Type Instruction prompt (ES) Instruction prompt (TR)
System
Instruct.
Tarea: Generar una sentencia Cypher para consultar una base de datos de grafos. Instrucciones: Usa únicamente los tipos de relaciones y propiedades proporcionados en el esquema. No utilices ningún otro tipo de relación ni propiedades que no estén incluidas en el esquema. No incluyas explicaciones ni disculpas en tus respuestas. No respondas a ninguna pregunta que no sea una solicitud para construir una sentencia Cypher. No incluyas ningún texto excepto la sentencia Cypher generada. Görev: Bir çizge veritabanını sorgulamak için bir Cypher ifadesi oluştur. Talimatlar: Yalnızca şemada verilen ilişki türlerini ve özellikleri kullan. Şemada verilmeyen herhangi bir ilişki türünü veya özelliği kullanma. Yanıtlarında hiçbir açıklama veya özür ifadesine yer verme. Cypher ifadesi oluşturmak dışında başka bir şey isteyen sorulara yanıt verme. Oluşturulan Cypher ifadesi dışında hiçbir metin ekleme.
User
Instruct.
Genera una sentencia Cypher para consultar una base de datos de grafos. Usa únicamente los tipos de relaciones y propiedades proporcionados en el esquema.
[…]
Bir çizge veritabanını sorgulamak için Cypher ifadesi oluştur. Şemada verilen ilişki türleri ve özellikler dışında hiçbir şeyi kullanma.
[…]
4.2. Impact of prompt language
We also translated the prompts into Spanish and Turkish to evaluate the impact of prompt language on model performance. The prompts were translated using the GPT-4o-mini model and are shown in Table 3. The performance comparison for Spanish and Turkish with and without translated prompts are presented in Figure 3. The results show that when prompts are translated, the ROUGE-L scores, which used for translation-based evaluation comparing ground-truth and generated text, improve slightly (by about 1–1.5%). However, Exact-Match scores for execution-based evaluation remain largely unchanged. These results suggest that prompt translation has a minor impact on the overall result in our setup. It is important to note that in our dataset, the database schemas, including nodes, relationships and properties, remain in English. Future work could explore the impact of localizing schema items to match the language of the input and prompt, which may reveal further interactions between language and performance.
(a) Translation-based evaluation: ROUGE-L score
(b) Execution-based evaluation: Exact-Match score
Figure 3. Model performance with/out prompt translation.
5. Conclusion
This work presents a performance comparison of foundational LLMs on the Text2Cypher task across English, Spanish, and Turkish. We created and released a multilingual test set by translating English questions into Spanish and Turkish while keeping the same ground-truth queries, enabling fair cross-lingual comparison of model performance. Our results consistently show that foundational models perform best on English questions, followed by Spanish, and lowest on Turkish. This pattern is likely driven by differences in training data availability and the linguistic characteristics, which pose additional challenges for the models. We also explored the impact of translating task instruction prompts alongside the input questions. Our findings suggest that prompt translation slightly improves translation-based evaluation but has little to no effect on query execution results, suggesting that prompt language has minimal impact on overall performance in our setting. Future work includes schema localization to improve multilingual performance, fine-tuning models on individual or combined languages and expanding evaluation to additional languages.